# SkyMask: Attack-agnostic Robust Federated Learning with Fine-grained Learnable Masks

### v1

@@ IV-B2 Under no attack
CIFAR10 (ResNet20) IID None 0.75 0.79 0.79 / / / / / /
LF 0.73 0.75 0.78 3.45% 0.00% 3.24% 5.80% 4.26% 0.00%
Min-Max 0.74 0.25 0.79 0.05% 100% 0.00% 0.00% 100% 0.00%
Min-Sum 0.74 0.31 0.79 0.14% 100% 0.00% 1.00% 100% 0.00%
Fang-Trim 0.74 0.77 0.78 0.03% 0.00% 0.00% 0.00% 0.05% 0.00%
Fang-Krum 0.74 0.18 0.79 4.04% 94.60% 0.00% 10.60% 100% 0.00%
Scaling 0.79 / 0.09 0.78 / 0.10 0.79 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
HAR (LR) None 0.96 0.92 0.96 / / / / / /
LF 0.96 0.68 0.96 0.22% 13.27% 0.00% 0.13% 100% 0.00%
Min-Max 0.96 0.07 0.96 1.38% 100% 0.83% 2.00% 100% 0.00%
Min-Sum 0.95 0.53 0.96 31.6% 100% 4.53% 32.0% 100% 4.00%
Fang-Trim 0.95 0.96 0.96 0.00% 0.00% 0.96% 0.00% 0.00% 0.00%
Fang-Krum 0.95 0.13 0.96 26.4% 100% 0.03% 29.3% 100% 0.00%
Scaling 0.96 / 0.20 0.95 / 0.19 0.96 / 0.20 0.25% 0.00% 0.00% 12.58% 0.00% 0.08%
IV-B3 Under a high fraction of attacks
The experiments mentioned in the previous sections set the fraction of malicious clients to 20%, while in a real application environment, an attacker can control a larger fraction of malicious clients for the attack. Therefore, to verify the defensive capability of our system in this case, we use CNN as the global model on Fashion-MNIST dataset in the non-IID case and conduct comparison experiments with 40%, 60%, and 80% of malicious clients. Since Bulyan requires that the number of malicious clients $m$ should satisfy $n\geq 4m+3$, these experiments do not include this algorithm. Trim requires that the number of malicious clients should meet $n\geq 2m+1$, and Multi-Krum requires that $n-c>2m+2$, so these aggregation algorithms are only included in 40% experiments.
The experimental results are shown in Figure 4. Most algorithms perform well under Scaling attack except for FedAvg and Krum. When the fraction of malicious clients is 40%, Trim loses its defensive capability completely, and Multi-Krum can only work under LF attack. Krum can reduce the impact of attacks under LF and Fang-Trim attacks, but the accuracy loss is still large. Although the accuracy of FLTrust fluctuates in some communication rounds, it still tends to converge to an unattacked level. From Figure 4(c), Figure 4(d) and Figure 4(e), we can see that SkyMask converges more stable than FLTrust when they are attacked by Fang-Krum attack, Min-Max attack and Min-Sum attack. In Figure 4(a) and Figure 4(e), the accuracy obtained by SkyMask is 2% to 4% higher than FLTrust’s under LF and Min-Sum attacks.
When the fraction is more than 50%, only FLTrust and SkyMask maintain the defensive capability, but FLTrust’s performance worsens as the fraction rises. FLTrust’s testing accuracy fluctuates even more, and the global model even converges in the wrong direction (e.g., in Figure 4(q), the accuracy obtained by FLTrust decreases from 85% to 81% after a sharp fluctuation). Under other attacks, the increasing fraction influences FLTrust’s convergence speed, and SkyMask converges faster than it.
(a) Fang-Krum attack, Fashion-MNIST
(b) Min-Sum attack, Fashion-MNIST
(c) LF attack, CIFAR-10
Figure 5: The impact of detection failure and the training efficiency of Tolpegin defense, FLDetector and SkyMask.
(a) On FashionMNIST Dataset (non-IID) with CNN
(b) On FashionMNIST Dataset (IID) with CNN
(c) On CIFAR-10 Dataset (non-IID) with ResNet20
(d) On CIFAR-10 Dataset (IID) with ResNet20
Figure 6: The impact of root data distribution on SkyMask.
IV-C The Significance of Learnable Masks
@@
We conduct comparative experiments on SkyMask and the other malicious client detection algorithms, i.e., Tolpegin defense [2] and FLDetector [16] (combined with Median). We sample the result of malicious client detection every ten communication rounds and calculate all the metrics described in §IV-A7.
In Table II, we can see that FLDetector only works under LF attack, Fang-Trim attack and Scaling attack. Min-Max and Min-Sum attacks make FLDetector completely confuse the malicious model updates with the benign model updates.
In the Fashion-MNIST part of Table II, we can see that it is hard for Tolpegin defense to detect malicious clients stably, especially under Fang-Krum attack and Min-Sum attack. Its FPR and FNR are pretty high, so it fails to defend against attacks effectively and loses a lot of benign information. Quite the opposite, SkyMask perfectly completes the detection task and achieves an FNR of less than 4‰  on Fashion-MNIST dataset with a CNN model, which means that the server hardly misses any malicious client, and it can eliminate the harm of attacks. SkyMask also achieves an FPR of less than 3‰ , so it can preserve most benign clients and better characterize the global data distribution.
Figure 5(a) and Figure 5(b) show the trend of testing accuracy on non-IID Fashion-MNIST with a CNN global model under Fang-Krum attack and Min-Sum attack. Detection failure means that the FL defense system does not detect all the malicious clients in that round. We can see that SkyMask gradually rises and converges without detection failure. On the contrary, Tolpegin defense converges but with poor testing accuracy under Fang-Krum attack, and it is damaged under Min-Sum attack. Each detection failure causes a fluctuation in testing accuracy.
@@
In the CIFAR-10 part of Table II, Tolpegin defense performs even worse in the non-IID case with large model architecture (ResNet20). Except for Fang-Trim attack, all other attacks greatly harm it. The large FNR of Tolpegin defense means many malicious clients complete their attacks. The large FPR of Tolpegin defense shows that the remaining benign clients may lose some representation, so Tolpegin defense achieves a lousy performance. In the IID case, Tolpegin defense resumes partial defensive capability, but there is still about 10% FNR under Fang-Krum attack. However, our SkyMask achieves the same prediction accuracy as a no-attack case, with FPR less than 5% and FNR less than 3%.
LF attack on CIFAR-10 dataset is more stealthy, and the FNR of SkyMask rises to 2.6%. As Figure 5(c) shows, at the beginning of FL training under LF attack, determining whether the difference between clients comes from data distribution or malicious client’s anomaly operation is more challenging. However, this does not matter since the LF attack is weak at the beginning of the training process. The global model still moves towards convergence; after several communication rounds, SkyMask can resume normal detection ability and identify these malicious clients. This mistake does not prevent our method from helping the global model reach convergence and high testing accuracy.
TABLE III: The impact of threshold in binarization on SkyMask.
Dataset (Model) Attack Acc. FPR FNR
0.3 0.4 0.5 0.6 0.7 0.3 0.4 0.5 0.6 0.7 0.3 0.4 0.5 0.6 0.7
Fashion -MNIST (CNN) non-IID None 0.89 0.89 0.89 0.89 0.89 / / / / / / / / / /
LF 0.88 0.88 0.89 0.89 0.88 0.01% 0.02% 0.00% 0.03% 0.32% 0.08% 0.02% 0.02% 0.00% 0.04%
Min-Max 0.89 0.89 0.89 0.89 0.89 0.00% 0.00% 0.27% 0.00% 0.00% 0.00% 0.00% 0.40% 0.00% 0.00%
Min-Sum 0.88 0.89 0.89 0.89 0.88 0.04% 0.20% 0.26% 0.34% 0.89% 0.00% 0.40% 0.00% 0.00% 1.60%
Fang-Trim 0.88 0.89 0.89 0.88 0.88 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.89 0.89 0.89 0.89 0.18% 0.28% 0.20% 0.40% 0.28% 0.40% 0.00% 0.00% 0.40% 0.00%
Scaling 0.89 / 0.10 0.89 / 0.12 0.89 / 0.10 0.89 / 0.10 0.89 / 0.11 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fashion -MNIST (CNN) IID None 0.88 0.89 0.90 0.89 0.89 / / / / / / / / / /
LF 0.89 0.89 0.90 0.90 0.89 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Max 0.89 0.90 0.90 0.90 0.89 0.08% 0.00% 0.00% 0.00% 0.01% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.89 0.89 0.90 0.90 0.89 0.04% 0.05% 0.01% 0.01% 0.34% 0.13% 0.00% 0.00% 0.00% 1.00%
Fang-Trim 0.89 0.90 0.90 0.90 0.90 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.89 0.90 0.90 0.89 0.77% 0.56% 0.19% 0.05% 0.07% 1.00% 0.20% 0.00% 0.00% 0.20%
Scaling 0.89 / 0.12 0.90 / 0.11 0.90 / 0.10 0.90 / 0.11 0.90 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
CIFAR-10 (ResNet20) non-IID None 0.76 0.76 0.76 0.76 0.76 / / / / / / / / / /
LF 0.76 0.76 0.76 0.76 0.76 4.72% 4.33% 2.03% 2.40% 6.65% 5.70% 2.80% 2.60% 4.00% 6.30%
Min-Max 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Trim 0.76 0.77 0.76 0.76 0.76 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.77 0.77 0.77 0.77 0.77 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Scaling 0.76 / 0.09 0.76 / 0.12 0.77 / 0.12 0.77 / 0.10 0.77 / 0.10 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
CIFAR-10 (ResNet20) IID None 0.79 0.78 0.79 0.78 0.78 / / / / / / / / / /
LF 0.76 0.78 0.78 0.79 0.78 5.31% 3.40% 3.24% 1.25% 8.04% 5.45% 1.75% 0.00% 4.45% 8.45%
Min-Max 0.78 0.79 0.79 0.79 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Sum 0.78 0.79 0.79 0.79 0.79 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Trim 0.78 0.79 0.78 0.78 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.78 0.79 0.79 0.78 0.78 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Scaling 0.79 / 0.11 0.79 / 0.10 0.79 / 0.10 0.79 / 0.09 0.79 / 0.11 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
HAR (LR) None 0.96 0.96 0.96 0.96 0.96 / / / / / / / / / /
LF 0.96 0.96 0.96 0.96 0.96 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Min-Max 0.96 0.96 0.96 0.96 0.96 1.70% 1.57% 0.97% 2.50% 0.43% 1.00% 0.00% 1.00% 1.00% 1.40%
Min-Sum 0.96 0.96 0.96 0.96 0.96 2.17% 2.63% 0.87% 3.60% 5.70% 2.50% 2.00% 0.70% 1.00% 3.00%
Fang-Trim 0.96 0.96 0.96 0.96 0.96 0.16% 0.00% 0.00% 0.00% 0.00% 2.63% 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.96 0.96 0.96 0.96 0.96 0.70% 0.60% 0.00% 0.07% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Scaling 0.96 / 0.22 0.96 / 0.19 0.96 / 0.20 0.96 / 0.16 0.96 / 0.20 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00%
Figure 7: The impact of root data distribution on model convergence.
For HAR dataset, Tolpegin defense’s testing accuracy seems good even when FNR rises to about 30%. The reason is that since the classification task is simple, even if detection failure in some rounds leads to the fluctuation of the testing accuracy, the testing accuracy can still rise to a reasonable level through successive rounds of detection success. In contrast, SkyMask achieves a good malicious client detection result with FNR less than 4%, maintaining high testing accuracy.
IV-D Sensitivity analysis
(a) On FashionMNIST Dataset (non-IID) with CNN
(b) On FashionMNIST Dataset (IID) with CNN
(c) On CIFAR-10 Dataset (non-IID) with ResNet20
(d) On CIFAR-10 Dataset (IID) with ResNet20
(e) On HAR Dataset with LR
Figure 8: The impact of dimensionality reduction on SkyMask.
IV-D1 The Impact of Root Dataset Distribution
The server’s root dataset is randomly sampled from the global data distribution by default. It is most likely to collect a root dataset with bias in the real world. However, we can prove that as soon as the bias is not too large ($p\leq 0.8$), our method can still take advantage of the root dataset and complete the malicious client detection task. We experiment on Fashion-MNIST with a CNN global model and CIFAR-10 with a Resnet20 model. The results are as testing accuracy part of Figure 6 shown: most of the testing results of the global model obtained by our method are very close or equal to the model obtained by FedAvg under no attack. However, in some cases, such as Figure 6(b) under Min-Max attack, and Figure 6(d) under no attack and LF attack, the testing accuracy is not acceptable when the bias probability of the root dataset is 0.95. Considering that the root dataset has a total of 100 samples, the lack of some classes of samples in the root dataset is possible when the bias probability of the root dataset is 0.95, so this result is reasonable.
The FPR and FNR results in the figures show that the malicious client detection accuracy of SkyMask fluctuates as the degree of non-IID of the root data changes but remains very low. Our method does not fail until the bias probability of the root dataset exceeds 80%. As long as the generated root dataset has a bias not so significant, SkyMask can complete its malicious client detection task and keep the global model a good performance.
In Figure 7, we show the variation curve of accuracy of SkyMask for non-IID FashionMNIST experiments. The convergence speed reduces along with the data distribution bias increasing, especially under Fang-Trim attack and Scaling attack. Although the root dataset’s different bias probability ($p\leq 0.8$) does not affect the convergence result, this phenomenon tells us that the imbalanced root dataset can slow down the global model’s convergence, and we had better choose a good root dataset for the server.
IV-D2 The Impact of Thresholds in Binarization
As the description of the processing of binary masks in Section III-C, the server should choose a threshold $\tau$. Our optimization strategy makes the parameters in masks converge to 0 or 1, so SkyMask should not be sensitive to threshold $\tau$. We conduct experiments with different thresholds from 0.3 to 0.7, and the results are shown in Table III.
On Fashion-MNIST dataset, it can be seen that SkyMask loses a little accuracy under some attacks when the threshold goes far away from 0.5. Still, it also has a similar accuracy as unattacked FedAvg’s level, with about 1% accuracy loss. SkyMask achieves FNR lower than 1.6%, so there are almost no undetected malicious clients; FPR is lower than 0.9%, so most benign clients are retained.
On CIFAR-10 dataset, especially under LF attack, FPR and FNR increase as the threshold deviates from 0.5. But they are still less than 9%, and the testing accuracy loss is less than 1%. And for HAR, we can see similar results that SkyMask keeps a very low FPR and FNR, achieving the same testing accuracy as unattacked FedAvg. Therefore, the change in threshold does not affect our system’s performance, and SkyMask is not sensitive to the threshold in binarization.
IV-D3 The Impact of Dimension Reduction
The clustering results of SkyMask are influenced by the dimensionality of the features after dimensionality reduction. In Figure 8, we can see that when the number of features after dimensionality reduction rises from 2 to 15, although the testing accuracy does not decrease severely in most cases, there is a noticeable increase in the FPR and FNR, especially under LF attack. The reason is that the framework uses the Gaussian mixture model to cluster the dimension-reduced masks. Since the total number of masks is not large enough, the clustering results may not be satisfying in high-dimension space due to the curse of dimensionality. Therefore, we set the number of features after dimensionality reduction as two empirically.
IV-E SkyMask’s Scalability
To demonstrate the scalability of SkyMask, we conducted experiments utilizing a CNN and ResNet20 global model trained on Fashion-MNIST and CIFAR-10, respectively. For Fashion-MNIST, we assessed the prediction accuracy of the main task, along with the FPR and FNR of SkyMask under various attack scenarios, encompassing 500 and 1000 clients. We have mentioned in §IV-A1 that the local datasets assigned to the clients were disjoint and each client’s dataset should be reasonably sized, so for the CIFAR-10 experiments with fewer data samples, we selected 200 and 500 clients.
Table IV shows that SkyMask can work well even if the number of clients increases to 1000 on Fashion-MNIST. As the number of clients increases, more malicious clients are missed under Min-Sum and Fang-Krum attack, but the FNR is still lower than 1.6%. In addition, the missed malicious local models do not affect the performance of the global model. We can see that the global model of SkyMask achieves the same testing accuracy as the unattacked global model’s obtained by FedAvg.
For the experiments on CIFAR-10, we observed a decrease in the testing accuracy results of the unattacked FedAvg as the total number of clients increased. This can be attributed to the challenges posed by the smaller size of the local datasets and the larger number of clients in the FL training process [23]. However, our SkyMask algorithm demonstrated robust performance even under most attacks. For the LF attack, SkyMask achieved comparable testing accuracy to the unattacked FedAvg, despite a slight increase in the FNR to approximately 9%.
Therefore, SkyMask can work well when the total number of clients increases. Even if there are millions of clients in the FL system, the server can randomly choose a fixed number (e.g., 100 or 500) of clients in each communication round to execute malicious client detection and aggregation. The details we will not describe in this paper.
TABLE IV: The impact of the total number of clients. (The first line in table is the testing accuracy result of FedAvg under no attack.)
FashionMNIST (CNN) Test Acc. FPR FNR
500 1000 500 1000 500 1000
FedAvg 0.89 0.89 / / / /
None 0.89 0.89 / / / /
LF 0.89 0.89 0.40% 0.40% 0.94% 0.62%
Min-Max 0.89 0.89 0.40% 0.30% 0.00% 0.00%
Min-Sum 0.89 0.89 0.40% 1.60% 0.00% 0.00%
Fang-Trim 0.89 0.89 0.00% 0.00% 0.00% 0.00%
Fang-Krum 0.89 0.89 0.60% 1.20% 0.00% 0.00%
Scaling 0.89 / 0.11 0.89 / 0.10 0.00% 0.00% 0.00% 0.00%
@@ IV-E SkyMask’s Scalability
V Discussion and Limitations
The current Byzantine attacks are launched with either a single client or multiple clients executing the same attack strategy, which may only account for an increased proportion of the model aggregation stage on the server, but the malicious clients do not collaborate. Increasingly, some researchers find that some specific layers are more vulnerable to Byzantine attack [24]. We also observe that some Byzantine attacks modify parameters near the output layer significantly more than those far from the output layer.
Therefore, each client can be responsible for a different part of model poisoning, using the philosophy of “not putting eggs in the same basket.” The attacker hides poisoning factors in all the uploaded models of the malicious clients it compromises. When most or all malicious model updates bypass the malicious client detection mechanism and get aggregated on the server, it can form a powerful attack on the global model, which is truly a collaborative attack by multiple malicious clients.
The damage to the characteristics of local model updates by this collaborative attack is unknown, and it remains to be verified whether SkyMask can efficiently detect these malicious attacks. For other malicious client detection algorithms, detecting this abnormal pattern in a single model update may be challenging. Theoretically, using our mask mechanism, the collaborative impact of all malicious clients on the global model could be shown after masked aggregation so that the optimization process can change their corresponding masks’ pattern. Actually, Fang attack and AGR-agnostic attack have attempted to attack collaboratively, which weakens the anomaly level of model updates uploaded by each malicious client but influences the performance of the global model when all attacks are combined. Hence, there is a chance for our SkyMask to detect this kind of malicious client.
In the future, we will continue to study this topic in-depth, first verify the effectiveness of this collaborative attack, and deepen the ability of our learnable binary mask to extract the features of model updates and realize the defense capability for this unknown attack. In addition to collaborative attack, the performance of SkyMask under multiple attack types in one communication round is unknown. We should explore how many clusters will the corresponding masks obtained by SkyMask form, and we should also adapt the settings of SkyMask according to this.
Furthermore, SkyMask can be seen as an additional module to the aggregation algorithm. Not only the basic FedAvg algorithm, the SOTA FL algorithms, such as FedProx [25], FedBN [26], or FedAMP [27], can also be equipped with our malicious client detection module. This feature lets the FL system retain the heterogeneity-handling, high convergence rate, or personalization features the original aggregation algorithm may have. It is also possible to enjoy the defensive capability of our malicious client detection strategy, and we will explore the merits of various combinations in future work.
VI Conclusion
@@
We propose a new attack-agnostic robust FL framework called SkyMask to defend against Byzantine attacks. By training parameter-level learnable binary masks on a clean root dataset, SkyMask is the first to implement a fine-grained detection of the poisoned elements of local model updates. Because SkyMask analyzes the mask as a whole, it avoids accidentally removing outliers caused by data heterogeneity on some elements of model updates and precisely removes the malicious clients. Extensive experiments on IID and non-IID data under different attacks prove the effectiveness of SkyMask. Our method shows a solid defensive capability, better than various robust aggregation algorithms and existing malicious client detection methods. Moreover, it also tackles the problem that most clients are malicious and can defend against attacks with a fraction of malicious clients up to $80\%$.

### v2

@@ 4.2 The Defense Effectiveness of SkyMask
DBA 0.65/0.47 0.77/0.10 0.77/0.10 0.00% 0.00% 0.00% 46.1% 0.00% 0.00%
4.3 The Significance of Learnable Masks
@@
We conduct comparative experiments on SkyMask and the other malicious client detection algorithms, i.e., Tolpegin defense [30] and FLDetector [36]. We sample the result of malicious client detection every ten communication rounds and calculate all the metrics described in §4.1.5.
@@
In Table 2, we can see that FLDetector only works under LF attack, Fang-Trim attack and targeted attack. Min-Max and Min-Sum attacks make FLDetector completely confuse the malicious model updates with the benign model updates. Tolpegin defense only survives under Fang-Trim attack, and all other attacks greatly harm it. The large FNR of Tolpegin defense means many malicious clients complete their attacks. The large FPR of Tolpegin defense shows that the remaining benign clients may lose some representation, so Tolpegin defense achieves a lousy performance. Our SkyMask achieves the same prediction accuracy as a no-attack case, with FPR less than 5% and FNR less than 3%.
Fig. LABEL:fig:detect_a and Fig. LABEL:fig:detect_b show the trend of testing accuracy on Fashion-MNIST with a four-layer CNN global model under Fang-Krum and Min-Sum attack. Detection failure means that the FL defense system does not detect all the malicious clients in that round. Under Fang-Krum attack, Tolpegin defense converges but yields poor testing accuracy, while FLDetector is rendered ineffective. Additionally, both FLDetector and Tolpegin defense exhibit fluctuating and low testing accuracy under Min-Sum attack due to detection failures. In contrast, SkyMask stably converges without encountering any detection failures.
Table 3: The impact of the total number of clients. (The first line in table is the testing accuracy result of FedAvg under no attack.)
@@
DBA 0.75/0.10 0.73 /0.10 0.00% 0.00% 0.00% 0.00%
4.4 SkyMask’s Scalability
To demonstrate the scalability of SkyMask, we conduct experiments utilizing a ResNet20 global model trained on CIFAR-10. We assess the prediction accuracy of the main task, along with the FPR and FNR of SkyMask under various attack scenarios, encompassing 200 and 500 clients.
In Table 3, we observe a decrease in the testing accuracy results of the unattacked FedAvg as the total number of clients increased. This can be attributed to the challenges posed by the smaller size of the local datasets and the larger number of clients in the FL training process [12]. For the LF attack, SkyMask achieves comparable testing accuracy to the unattacked FedAvg, despite a slight increase in the FNR to approximately 9%. For all fine-grained attacks and targeted attacks, SkyMask maintains FPR and FNR at 0%. Our SkyMask algorithm demonstrates robust performance under all attacks even when the total number of clients significantly increases.
5 Conclusion
@@
We propose a new attack-agnostic robust FL framework called SkyMask to defend against Byzantine attacks. By training parameter-level learnable binary masks on a clean root dataset, SkyMask is the first to design a fine-grained detection of the poisoned elements of local model updates. Extensive experiments on non-IID datasets under different attacks prove the effectiveness, generality, and scalability of SkyMask. Our SkyMask shows a solid defensive capability, better than various robust aggregation algorithms and existing malicious client detection methods. Moreover, our SkyMask also tackles the problem that most clients are malicious and can defend against attacks with a fraction of malicious clients up to $80\%$.
Acknowledgements
This work was partially supported by the National Key R&D Program of China (2022YFB4402102), the Shanghai Key Laboratory of Scalable Computing and Systems, the HighTech Support Program from STCSM (No.22511106200), and Intel Corporation (UFunding 12679). Tao Song is the corresponding author.